OpenStudy (anonymous):

If (8z − 9)(8z + 9) = az2 − b, what is the value of a? The value of a is ____________.

OpenStudy (anonymous):

Distribute (8z − 9)(8z + 9) on the left side of the equation first.

OpenStudy (anonymous):

64z+81=az2-b?

OpenStudy (anonymous):

Yes, so 64 is a.

OpenStudy (anonymous):

Thank you so much!!

OpenStudy (anonymous):

The equation is actually 64z^2-81=az^2-b, also can be (8z)^2−81. And, you're welcome!